# Impact pressure enhancement upon disk impact on a boiling liquid

### v1

@@ Abstract
We then compute $U_{\text{in}}/C_{\text{sound,L}}$ using values of the temperature-dependent properties in Eq. (6) from [26, 22, 24], see Supplementary Material [27]. The fitting parameter $\lambda$ is taken to be $2.1$ as obtained by fitting the experimental $h_{0}$ data (not shown here). Replotting the rescaled maximum impact pressure at the disk center from Fig. 2 against the computed $U_{\text{in}}/C_{\text{sound,L}}$ gives a decent collapse of data as shown in Fig. 4, suggesting that this simple model predicts the dependence of the vapor condensation in the pocket on the impact velocity and ambient temperature during boiling liquid impact. In addition this confirms that the condensation-induced collapse of the vapor pocket is responsible for the very high impact pressures exerted on the impacting surface.
@@
The mechanism by which condensation creates the anomalously high maximum central impact pressures is the following: When phase change does not play a role, the solid-liquid impact is cushioned by the (relatively slow) compression of the gas/vapor pocket. If the vapor condenses sufficiently fast, the liquid reaches the solid liquid interface unbraked, resulting in a large and sudden impact pressure peak.

### v2

@@ Abstract
We then compute $U_{\text{sp}}/C_{\text{sound,L}}$ using values of the temperature-dependent properties in Eq. (6) from [aminian2022ideal, widiatmo2001equations, rausch2015density], see Supplementary Material [supplemental]. The fitting parameter $\lambda$ is taken to be $2.1$ as obtained by fitting the experimentally measured $h_{0}$ (not shown here, but see [fan2025disk]). Replotting the rescaled maximum impact pressure at the disk center from Fig. 2 against the computed $U_{\text{sp}}/C_{\text{sound,L}}$ gives a decent collapse of data as shown in Fig. 4, suggesting that this simple model predicts the dependence of the vapor condensation in the pocket on the impact velocity and ambient temperature during boiling liquid impact. In addition, this confirms that the condensation-induced collapse of the vapor pocket is responsible for the very high impact pressures exerted on the impacting surface.
@@
The mechanism by which condensation creates the anomalously high maximum central impact pressures is the following: When phase change does not play a dominant role, the solid-liquid impact is cushioned by the (relatively slow) pressurization of the vapor pocket. If the vapor condenses sufficiently fast, the liquid reaches the solid interface unbraked, resulting in a large and sudden impact pressure peak.
@@
Data availability: The data that support the findings of this article are openly available [dataSet].